what is the smallest sample size needed to estimate the mean number of days lost due to workplace injuries with error no more than 1 day and 95% confidence, given that the standard deviation is 0.5?
Answer by ewatrrr(24785)

 
Hi
what is the smallest sample size needed to estimate the mean number of days
lost due to workplace injuries with error no more than 1 day and 95%
confidence, given that the standard deviation is 0.5?
95% confidence ...z = 1.96
E = z*SD/sqrt(n)
n = (z/E *SD)^2
n = (1.96/1*.5)^2 = 15.3664
Smallest size would be 16
See Summary of vaious z values for different confidence levels
a a/2 crtical regions
90% 0.1 5% z <-1.645 z >+1.645
95% 0.05 2.50% z <-1.96 z >+1.96
98% 0.02 1% z <-2.33 z >+2.33
99% 0.01 0.50% z<-2.575 z >+2.575